Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 159     Entry time: 08 Sep 2024, 20:30
Author: Peter Zalden 
Group: FXE 
Subject: Issue 

We observe a continuous horizontal beam drift in SA1 after the M1 incident. It seems not to converge, so one can wonder where this will go....

Attached screenshot shows the M2 piezo actuator RY that keeps the beam at the same position on the M3 PBLM. These are about 130 m apart, The drift is 1V/12 hours. 0.01V corresponds to a beam motion by 35 um. So the drift velocity is 0.3 mm/hour. In the past days this had at least once lead to strong leakage around M2. But the two feedbacks in the FXE path are keeping the beam usable most of the time.

Attachment 1: 2024-09-08-194936_755701995_exflqr51474.png  24 kB  Uploaded 08 Sep 2024, 21:37  | Hide | Hide all
2024-09-08-194936_755701995_exflqr51474.png
ELOG V3.1.4-7c3fd00